Question Help

InstituteInstitute

Corn Hole is a small business that

JacksonJackson

Morris developed while in college. He began building wooden corn hole game sets for​ friends, hand painted with college colors and logos. As demand​ grew, he hired some workers and began to manage the operation.

InstituteInstitute

Corn hole maintains two​ departments: contstruction and painting. In the construction​ department, the games require wood and labor.

InstituteInstitute

Corn Hole has some employees who have been with the company for a very long time and others who are new and inexperienced.

Requirement 1. For the construction​ department, calculate the price and efficiency variances for the wood and the price and efficiency variances for direct manufacturing labor. Label each variance as favorable​ (F) or unfavorable​ (U).

Begin by calculating the price and efficiency variances for the wood.

Price variance

Efficiency variance

Calculate the price and efficiency variances for direct manufacturing labor.

Price variance

Efficiency variance

Requirement 2. Record the journal entries for the variances incurred. ​(Record debits​ first, then credits. Exclude explanations from any journal​ entries.)

Prepare the journal entry for the direct materials price variance.

Journal Entry

Date

Accounts

Debit

Credit

Next prepare the journal entry for direct materials efficiency variance.

Journal Entry

Date

Accounts

Debit

Credit

Now prepare the journal entry for direct manufacturing labor price and efficiency variances.

Journal Entry

Date

Accounts

Debit

Credit

Requirement 3. Discuss logical explanations for the combination of variances that

InstituteInstitute

experienced.

JacksonJackson

paid

less

more

for the wood. As a​ result, the wood was

higher

lower

quality and workers used

less

more

of it.

JacksonJackson

used

less

more

experienced workers in September than he usually does. This resulted in payment of

higher

lower

wages per hour. The new workers were

less

more

efficient and took

fewer

more

hours than normal.

​Overall,

JacksonJackson​'s

total wage bill was

higher

lower

than expected.

InstituteInstitute

Corn Hole uses standard costing for the game sets.

JacksonJackson

expects that a typical set should take

4

hours of labor in the construction​ department, and the standard wage rate is

$10.00

per hour. An average set uses

22

square feet of​wood, allowing for a certain amount of scrap. Because of the nature of the​ wood, workers must work around the flaws in the materials.

JacksonJackson

shops around for good deals and expects to pay

$4.30

per square foot.

JacksonJackson

does not store​ inventory, and buys the wood as he receives an order.

For the month of​ September,

JacksonJackson​'s

workers produced

7575

corn hole sets using

308

hours and

1,750

square feet of wood.

JacksonJackson

bought wood for $7,200

​(and used the entire​ quantity), and incurred labor costs of$2,945.

1.

For the construction​ department, calculate the price and efficiency variances for the wood and the price and efficiency variances for direct manufacturing labor.

2.

Record the journal entries for the variances incurred.

3.

Discuss logical explanations for the combination of variances that the construction department of

InstituteInstitute

Corn Hole experienced.
Answer #1

Answer 1:

Price and efficiency variances for the wood:

Standard price = $4.30

Standard quantity = 22 * 75 = 1650 square feet

Actual quantity = 1750 square feet

Actual cost = $7,200

Price variance = AP * AQ - SP * AQ = 7200 - 4.30 * 1750 = $325 Favorable

Efficiency variance = AQ * SP - SQ * SP = 1750 * 4.30 -1650 * 4.30 = $430 Unfavorable

$325 Favourable $430 Unfavorable Price variance Efficiency variance

Price and efficiency variances for direct manufacturing labor:

Standard wage rate (SR) = $10.00

Standard labor hours (SH) = 4 * 75 = 300 Labor Hours

Actual labor hours (AH) = 308

Actual labor costs = $2,945.

Price variance = AR * AH - SR * AH = 2945 - 10 * 308 = $135 Favorable

Efficiency variance = AH * SR - SH * SR = 308 * 10 - 300 * 10 = $80 Unfavorable

$135 Favourable $80 Unfavorable Price variance Efficiency variance

Answer 2:

Journal entry for the direct materials price variance:

Journal Entry Accounts Debit Credit Date Inventory $7,525 Sept. $325 Mterial Price Variance Accounts Payable Cash $7,200

Journal entry for direct materials efficiency variance:

Journal Entry Debit Credit Date Accounts Work In process $7,095 $430 Sept. Material Efficiency Variance $7,525 Inventory

Journal entry for direct manufacturing labor price and efficiency variances:

Journal Entry Debit Credit Date Accounts Work In process $3,000 Sept. Labor Efficiency Variance $80 Labor Price/Rate Variance

Answer 3:

(a) Jackson Jackson paid less for the wood. As a​ result, the wood was lower quality and workers used more of it.

Explanation:

Price variance is favorable. Price per square feet paid is (7200 /1750=) $4.11 per square feet which is less than standard price of $4.30. Price variance is favorable but efficiency variance is unfavorable. This implies wood purchased is lower quality (purchased at lesser price) and workers used more of it.

(b) Jackson Jackson used less experienced workers in September than he usually does. This resulted in payment of lower wages per hour. The new workers were less efficient and took more hours than normal. Overall, Jackson Jackson​'s total wage bill was lower than expected.

Explanation:

Total standard labor cost = 75 * 4 * 10 = $3000

Actual cost = $2945

As such Jackson Jackson​'s total wage bill was lower than expected.

Price variance is favorable since Jackson Jackson used less experienced workers. Efficiency variance is unfavorable since less efficient workers took more time to complete.